### v1

# Training a Huggingface Model on AWS Sagemaker (Without Tears)

Abstract
Developments of Large language models (LLMs) have been mainly led by resource-rich research groups and industry partners. Without the luxury of on-premise computing resources to train LLMs of growing complexity, researchers are adopting cloud machine learning services like AWS Sagemaker to train Huggingface models. While a competent researcher can train a model well on a local machine, AWS Sagemaker’s learning curve might be more challenging for non-seasoned cloud users. As a result, researchers end up scouring the web for information to fix issues and knowledge gaps in existing documentation to use Huggingface models in AWS Sagemaker. In this demo paper, to bring broader democratization of cloud adoption, we put the information to a single point such that research can start from zero to train their first Huggingface model.
Training a Huggingface Model on AWS Sagemaker (Without Tears)
Liling Tan
1 Introduction
State-of-the-art large language models (LLMs) have grown in size and training difficulty (scao2022bloom; le-scao-etal-2022-language; smith2022using; alpaca). While 90% of the global IT spending is still on-premises and yet to migrate to the cloud (andyjassy2023), many research teams without heavy funding support do not have the luxury of on-premise computing cluster to train an LLM from scratch nor fine-tune a competitive state-of-the-art model with physical GPU machines.
Unlike bare-metal cloud instances, products like Amazon Web Services (AWS) Sagemaker, Google Colaboratory, and Microsoft Azure ML Studio provide fuss-free access to Jupyter notebooks on managed virtual servers. These “Jupyter as a Service" products abstract the users away from infrastructure setup and let researchers focus on data and model improvements with minimal Infra-/DevOps (Infrastructure or Development Operations) knowledge (10.1145/3093338.3104159; perkel2018jupyter).
Machine Learning Operationalization (MLOps) is the process of bringing machine learning models into production (app11198861; huyen2022designing; 10081336). An end-to-end MLOps framework on the cloud would start from research and development tools support like “Jupyter as a service" and the ability to measure, monitor and maintain the models created from researchers’ notebooks in an automated manner.
Figure 1: Sample code to train a Huggingface model using AWS Sagemaker
While there are a lot of books, documentation, blog posts, and a canonical online course [^1] https://huggingface.co/docs/sagemaker to educate users on HuggingFace with AWS Sagemaker, there’s a lack of a comprehensive guide to training a HuggingFace model for new cloud users. Figure 1 shows the options on the Huggingface model card pages to create sample code to train a model using AWS Sagemaker. The sample code is helpful as a start, but researchers soon find that the access to features in Sagemaker is different from how one would code the same model on a local machine.
This demo paper closes the knowledge gap between training a model on a local machine’s Jupyter Notebook and training on AWS Sagemaker. We will present a demo to train a HuggingFace model using Huggingface on AWS Sagemaker.
2 AWS Sagemaker
AWS Sagemaker is a managed cloud specifically used for machine learning. It comes with a suite of tools that ranges from data processing, Jupyter Notebook hosting, model provenance management, model training, hyperparameter tuning, and model inference hosting. In this demo, we will focus on model training and hyperparameter tuning and briefly touch on model deployment for the models we’ve trained in AWS Sagemaker.
2.1 Amazon SageMaker > Domains > Domain: mtsdemo > Jupyter Notebook
To begin, a researcher would navigate to the AWS Sagemaker tool on the search bar in the AWS console. However, unlike tools like Google Colab and Kaggle Notebook, the user is not brought directly into a Jupyter Notebook but is required to Setup a Sagemaker Domain to configure permissions and security controls.
For simplicity, a new user should use the quick setup; it takes 5-10 mins for the domain to be ready after the setup, after which it can be accessed from the left panel of the console through the Domains menu.
After clicking on the domain name, e.g., ‘mtsdomain‘, there will be an option to open a Studio page. And that will bring the user to Sagemaker Studio’s launcher page similar to Jupyter Lab on a local machine.
Next, on the top bar of the page, the user will need to create a new Jupyter notebook file using the File > New > Notebook and that will pop-up a Set up notebook environment window and user should configure the (i) image, (iii) instance type and (ii) kernel settings that fit the needs accordingly.
  • Image: This is the Docker image from which you want to load your notebook. [^2] If unsure, choose the latest CPU-optimized PyTorch and Python version compatible with the Huggingface transformer latest major/stable version
  • Instance: This is the instance type that you would like your Jupyter notebook to be hosted on. Note that this is not the instance type that you will train your model on in this demo, so a CPU instance works well. [^3] If unsure, choose ml.t3.medium, the comprehensive list of instance can be found on https://aws.amazon.com/ec2/instance-types/
  • Kernel: It is always Python since we work with Huggingface libraries.
Finally, after setting up the notebook environment, users will see the familiar Jupyter notebook page as one would experience on a local machine, Google Colab, or Kaggle notebook.
Figure 2: Shutting down virtual machine instances hosting the Jupyter notebook(s)
After the Jupyter notebook runs, the charges to host the ml.t3.medium instance will be initiated, even if the notebook is not running anything. Stopping the notebook functions or ending the Jupyter runtime does not stop the charges since the virtual machine [^4] For experienced cloud users, AWS calls them “Amazon Elastic Compute Cloud Instances”, Microsoft Azure calls them “Azure Virtual Machines”, Google Cloud Platform (GCP) calls them “Compute Engine: Virtual Machines”. We will refer to virtual hardware that AWS Sagemaker uses as “instances” and “virtual machine” interchangeably is still running and hosting notebook.
To stop the charges, frugal users are advised to open on the instance, kernel, apps, and terminal session control panel on the left bar, then click on the power-off button in the RUNNING INSTANCES and a pop-up will warn the user that shutting down the instances will stop all notebooks running on that instance and all unsaved data will be lost. [^5] Note that there can be multiple notebook kernels running on the same instance, esp. when you select the same instance type when starting a new notebook. As a good practice, turn off the KERNEL SESSIONS you do not need before shutting down the RUNNING INSTANCES or RUNNING APPS
The instructions in this section might seem trivial to seasoned AWS and Sagemaker users, but it can be daunting to new users and students learning to use cloud resources. While the instructions and user interface (UI) will change over time, the concepts on Jupyter services in MLOps remain; there will be some (i) hardware (virtual machine) hosting the notebook, i.e., the compute resources needed to run a Jupyter notebook and (ii) software (Docker image), i.e., the operating system and libraries that we will use to run the Jupyter notebook.
In the case of AWS Sagemaker, the controls of the (i) and (ii) falls on the user’s configurations, but with tools like Google Colab and Kaggle, they assume some defaults that most users will use and start with fixed (i) and (ii) configurations.
2.2 Amazon SageMaker > Training Jobs > Model Training
After setting the Sagemaker Studio and having a Jupyter notebook running in the studio, users would expect the Huggingface example from Figure 1 to work out of the box using the HuggingFace estimator. But, there is one more hardware setup before model training.
Unlike training a model on a local machine, where the machine running the Jupyter notebook is usually the same as the machine used to train the model, the instances that Sagemaker uses to run the Jupyter notebooks are generally not the machines that are used to train a model. One can think of the instance hosting the Jupyter Notebook as a controller machine that will turn on another machine or cluster of machines to start training the model, so the hosting for the notebook and the model training has different settings.
The number of instances permitted to host Jupyter Notebook and run model training are limited as per the default quota set by AWS. For first-time Sagemaker users, you would first need to request a quota increase to use a GPU instance for the model training.
User has to navigate to the Service Quotas > AWS services > Amazon SageMaker page, then type in Find quotas search bar the instance type you want to train on, e.g., ml.g4dn.2xlarge and select the ml.g4dn.2xlarge for training job usage option to enable the Request Quota Increase button, follow through the quota increase request pop-up and increase the quota as desired. [^6] If unsure, set the quota to 4; that means you can have a maximum of 4 concurrent ml.g4dn.2xlarge instance when training, users can use multiple instances to train a model or use multiple instances for a parallel hyperparameter grid search., [^7] Users can refer to Accelerated Computing section on https://aws.amazon.com/ec2/instance-types/ to find the appropriate instance type; for references, the P3 and P4 instances come Nvidia V100 and A100 GPUs and G4dn instance comes with T4 GPUs
It will take a couple of hours or days to see the updated quota before users can train a model with the Sagemaker HuggingFace estimator. [^8] Tip: From experience, the fastest approval for GPU instances comes from requesting only 1 ml.g4dn.xlarge instance for training job usage. The basic quota request on the platform through the automated Service Request page does not incur any charges, but to expedite or escalate the quota request with a tech support personnel, users can subscribe to various tiers of AWS Support Plans at additional cost. Users can track the status of the quota request from the Service Quotas > Quota request history on the left panel of the page.
2.2.1 Your First Huggingface Script on AWS Sagemaker
After getting the quota request approved, users can use the sample code (with slight modification) in Figure 1 to train their first Huggingface model on AWS Sagemaker. However, the verbatim sample code will not work out of the box since some task and model-specific hyperparameters need to be initialized.
Figure LABEL:working-sample (in Appendix) shows a working script adapted from the Huggingface sample code to launch a Python script that uses the transformers library with the sagemaker.huggingface.HuggingFace model estimator.
While the script is running, users can navigate to the Amazon SageMaker > Training > Training jobs page on the left panel and track the progress of the model training. For example, Figure 3 shows the dashboard of the successful/failed runs of the scripts that were launched using the sagemaker.huggingface.HuggingFace estimators as presented in Figure LABEL:working-sample.
Figure 3: Monitoring Model Trainings on the AWS Sagemaker Console
Up till this point, pointers and procedures presented in the above sections were unavailable in a single point of documentation; bits and pieces lurk around open tutorials, demonstrations and online courses. After a user manages to get a script shown in Figure LABEL:working-sample working, there are many other open and accessible demonstrations to use the various Huggingface libraries with AWS Sagemaker. Notably, we refer readers to start with https://huggingface.co/docs/sagemaker/train for further reading on the generic use of Huggingface models with AWS Sagemaker.
While documentation to fine-tune transformers models with preset datasets available on Huggingface datasets library are widely available, users are frequently asking ‘How do I train Huggingface model X with custom data Y with AWS Sagemaker?’ questions on the official Sagemaker section of the Huggingface community forum and StackOverflow’s huggingface-transformers and aws-sagemaker tagged questions. [^9] https://stackoverflow.com/questions/tagged/huggingface-transformers%20amazon-sagemaker, [^10] https://discuss.huggingface.co/c/sagemaker/17
For the rest of this demo paper, we will explain the sample code from Figure LABEL:working-sample and walk through the code line-by-line with the goal of empowering users to understand what they mean and how they should modify them to their specific NLP needs. As we explain the script, hopefully, it will answer Frequently Asked Questions (FAQs) on using Huggingface libraries with AWS Sagemaker. We will also highlight lesser-known AWS Sagemaker features to maximize developers’ productivity by emulating a researchers’ process to use a Jupyter Notebook on non-managed bare-metal virtual machines.
3 HuggingFace Estimator
We begin the script walkthrough starting from the HuggingFace estimator:
git_config = {
’repo’: ’https://github.com/’
’huggingface/transformers.git’,
’branch’: ’v4.36.0’}
huggingface_estimator = HuggingFace(
# Code location arguments.
git_config=git_config,
entry_point=’run_translation.py’,
source_dir=’./examples/’
’pytorch/translation’,
# AWS instance arguments.
role=role,
instance_type=’ml.g4dn.xlarge’,
instance_count=1,
# Docker image arguments.
transformers_version=’4.36.0’,
pytorch_version=’2.1.0’,
py_version=’py310’,
# Model’s hyperparameters.
hyperparameters = hyperparameters,
Other than the model’s hyperparameters that are dependent on how users coded the argument parser in their training script, there are three groups of Sagemaker-specific keyword arguments that the HuggingFace estimator expects (i) code location, (ii) AWS instance and (iii) Docker image arguments.
3.1 Specifying Code Location for HuggingFace Estimator
Looking at the git_config variable initialized before the HuggingFace estimator, we see that it is a Python dictionary pointing to the public GitHub repository https://github.com/huggingface/transformers.git and directing the estimator to pull the code from the v4.36.0 branch on the git tree.
The git_config together the source_dir and entry_point is asking the HuggingFace estimator to pull the run_translation.py script.
Q: Pulling from GitHub is nice, but how do I use my own Python script?
To train from a local script, specify the path to the local directory containing your .py file. For example, if you have a script/my_run_translation.py file in the same directory as Jupyter notebook that you are running the HuggingFace estimator, e.g., train_hf_sagemaker.ipynb,
/script
my_run_translation.py
train_hf_sagemaker.ipynb
Then users can remove the git_config variable and use the source_dir and entry_point arguments as follows:
huggingface_estimator = HuggingFace(
# Code location arguments.
entry_point=’my_run_translation.py’,
source_dir=’./script’,
# AWS instance arguments.
role=role,
instance_type=’ml.g4dn.xlarge’,
instance_count=1,
# Docker image arguments.
transformers_version=’4.36.0’,
pytorch_version=’2.1.0’,
py_version=’py310’,
# Model’s hyperparameters.
hyperparameters = hyperparameters,
There are other modes of utilizing the git_config argument of the estimator object; we refer readers to engdahl2008 comprehensive tutorial to integrate other GitHub repository or AWS CodeCommit.
Q: What if my custom training script uses libraries that are not supported by Huggingface transformers natively?
The simplest solution is a somewhat hidden AWS Sagemaker feature where you can enforce a SAGEMAKER_ENVIRONMENT variable that contains the requirements.txt file that one would usually use with Python pip in the same directory as the source_dir. For example, with a directory structure as follows,
/script
my_run_translation.py
requirements.txt
train_hf_sagemaker.ipynb
Users can specify the env argument, and Sagemaker will install the Python packages listed in the requirements.txt after starting up the virtual machine.
huggingface_estimator = HuggingFace(
# Code location arguments.
entry_point=’my_run_translation.py’,
source_dir=’./script’,
env={’SAGEMAKER_ENVIRONMENT’:
’requirements.txt’},
# AWS instance arguments.
role=role,
instance_type=’ml.g4dn.xlarge’,
instance_count=1,
# Docker image arguments.
transformers_version=’4.36.0’,
pytorch_version=’2.1.0’,
py_version=’py310’,
# Model’s hyperparameters.
hyperparameters = hyperparameters,
The snippets above with the fully functional code on Figure LABEL:working-sample (in Appendix) is sufficient to run an actual model training. There are more tips and tricks on Controlling AWS Instance for HuggingFace Estimator and Hyperparameter tuning jobs with Sagemaker in Appendix B and C.
4 Come for the Code, Stay for the Community
Open source and open knowledge documentation are fueled by the community. NLP researchers and ML practitioners are pragmatic, but for any open source to be successful, people come for the code and stay in the community (naomiclsc; drupalleander). The content presented in this demo is only possible with the bits and pieces of knowledge that many developers and researchers have recorded from the various forums.
As a final note to the readers, we want to encourage everyone to openly discuss and post and/or answer questions on the canonical forum for the Huggingface x AWS Sagemaker topic, https://discuss.huggingface.co/c/sagemaker/17. We hope to see the knowledge shared here improves with similar open documentation efforts.
5 Related Work
Unlike a typical demonstration paper, we are not presenting an NLP tool with a results table we can easily compare with other related work. Thus, we highlight several works introducing NLP tooling paradigms as further references to related NLP and MLOps literature.
kim-etal-2021-changes briefly described a no-code or low-code NLP experimentation setup to tune pre-trained models through an API without explicitly coding up the tuning procedures. While there seems to be little scientific contribution for researchers using AutoML, we concur with kim-etal-2021-changes on how low-/no-code setups are useful first-phase proof-of-concept tools before investing time and resources into coding up a script and running it on a Jupyter-as-a-service on the cloud.
Another line of work describes the data-centric MLOps contributions, notably, lhoest-etal-2021-datasets introduced the HuggingFace datasets hub that allows researchers to contribute, share and read datasets; their focus on API standardization, ease-of-use and community-drive data statement documentation (bender-friedman-2018-data) has springboard a new generation of open datasets sharing and allows open-ended discussion on biased or harmful data artifacts ramponi-tonelli-2022-features.
En vogue, Jupyter is the tool of choice to write NLP code and interact with the models (perkel2018jupyter). The closest comparable resources to our paper on model training models beyond Jupyter notebook are beyondjupyter introduction to environment setup and the HuggingFace Transformers with Sagemaker tutorial. However, they did not cover the various issues and resolutions that users will face when new to AWS Sagemaker.
6 Conclusion
We have demonstrated various tips and tricks to bridge the knowledge gaps using the Huggingface estimators on the AWS Sagemaker cloud service. Bringing the information from different blog posts, tutorials, and QnAs sites into this demo paper, we hope readers have a comprehensive guide to kick-start their Huggingface x AWS Sagemaker usage. Although this demo paper is not a traditional format of presenting a tool concerning results and tools, we introduced the engineering concepts in MLOps and practical steps to apply them in training NLP models.
Appendix
All the code snippets and Frequently Asked Questions listed on this paper will be hosted on https://github.com/anonymous/hfxaws-no-tears.
Like many NLP tools and deep learning library, the need to update the dependencies and code-base can be a never-ending pain. As the paper revision just before submitting the first draft to the conference, the minor versions of Huggingface transformers have changed at a monthly pace. We recommend the version controls with the solutions in Section 3.1 and 3.3.
Appendix A Limitations and Risk
Limitations (Engineering): This paper presented various knowledge gaps in training a Huggingface model on AWS Sagemaker. Although the content is hyper-limited to the Amazon Web Services cloud usage, the general MLOps tips presented in the form of AWS-related services are portable to other cloud providers (like Google Cloud Platform or Microsoft Azure) with their respective equivalent configurations. We refer readers to this cloudmap that maps the cloud services across the providers.
Limitations (Scientific): Although this paper presented the procedures and information to launch a Huggingface model training job on AWS Sagemaker, we have largely skipped the description on how to code entry_point script to build state-of-the-art NLP models. While this is in scope for the topics discussed in this paper, the code in the entry_point script is the same as how a researcher would train a model on a local machine. We acknowledge that there are plenty of good resources with little knowledge gaps on that, esp. on the cannonical Huggingface documentation site.
Risk: While there is no evident risk in the demonstrations presented in this paper. There are various security risks that cloud users should be concerned about. Using a self-contained on-premise environment, data stays within the organization’s server. But with cloud access, there will always be certain risks to cloud breaches (sabin2023; heiligenstein2023; thehackernews2023); we advise users to consult their organizations’ security administrators before sending private data onto the cloud.
Appendix B Controlling AWS Instance for HuggingFace Estimator
As discussed in Section 2.2, the quota issue will be the first blocker most users will face when setting the instance count. After increasing the quota and having the instance made available to Sagemaker for training jobs, users will soon find that only specific instance types are allowed to run the HuggingFace estimator.
Q: Aren’t the instance_type= and instance_count just boilerplate arguments?
Yes, they are; if you have finished training your model in a reasonable time and everything works without GPU memory limitation issues.
Suppose you encounter the cursed CUDA Out of Memory error, or the model is taking unreasonably long to converge. In that case, it is time to consider changing the AWS instances arguments to speed up or scale up your training jobs.
Q: How do I speed up or scale up my Huggingface model training with AWS Sagemaker?
For beginner users, we suggest following the simple Data Parallelism as suggested on multi-GPUs instances, e.g., ml.g4dn.2xlarge.
# configuration for running training
# on smdistributed data parallel
distribution = {
’smdistributed’:{
’dataparallel’:{
’enabled’: True }
huggingface_estimator = HuggingFace(
# Code location arguments.
entry_point=’my_run_translation.py’,
source_dir=’./script’,
# AWS instance arguments.
role=role,
instance_type=’ml.g4dn.2xlarge’,
instance_count=1,
# Docker image arguments.
transformers_version=’4.36.0’,
pytorch_version=’2.1.0’,
py_version=’py310’,
# Model’s hyperparameters.
hyperparameters = hyperparameters,
distribution = distribution
And when the model exceeds the computes possible on a single instance, then user can consider other parallelism options native to AWS or third-party parallelism libraries integrated into Huggingface transformers.
The AWS documentation page keeps an updated list of FAQs about picking AWS instances configurations to speed up training and/or reduce cloud spending budgets. We refer users to these tutorials to optimize AWS instance controls using HuggingFace estimators.
B.1 Selecting Docker Image for HuggingFace Estimator
As per writing this demo paper, the sample Sagemaker training codes from the Huggingface model cards have been referring to the de facto Huggingface v4.36.0 and PyTorch v2.1.0.
One of the issues that many users will encounter when setting the library versions to transformers<=4.36.0 is that the model they want to tune/train on is created post v4.36.0, e.g., users were excited to start using the Open AI Whisper model, but it is only supported in later versions of the transformers library not supported by then latest Sagemaker Docker image available.
It takes a while for new Sagemaker users to figure out which versions are supported in the preset docker images from the Deep Learning Containers (DLC) before they find links that navigate to the canonical list of supported DLC.
While the decision to make users explicitly state the transformers, pytorch, and python versions seems tedious instead of automatically fetching the latest image, it is a readability trade-off following the ‘explicit is better than implicit’ Pythonic satori.
Q: Can I Bring-Your-Own-Container (BYOC) to run the HuggingFace Estimator?
Yes, you can. If the account you are accessing the Sagemaker notebook from has access to the docker image or if you are using some publicly available docker image, you can use the image_uri instead of specifying the transformers, pytorch and Python versions.
ecr_uri=str(
"123456789012.dkr.ecr.eu-west-1."
"azmazonaws.com"
"/my-awsome-huggingface-docker:latest"
huggingface_estimator = HuggingFace(
# Code location arguments.
entry_point=’my_run_translation.py’,
source_dir=’./script’,
# AWS instance arguments.
role=role,
instance_type=’ml.g4dn.2xlarge’,
instance_count=1,
# Docker image arguments.
image_uri=ecr_uri,
# Model’s hyperparameters.
hyperparameters = hyperparameters
hotz2022 provided a detailed tutorial on modifying existing DLC images and modifying it before using them as a custom BYOC setup in AWS Sagemaker.
Appendix C Amazon SageMaker > Hyperparameter tuning jobs
While many tutorials on using Sagemaker stop at instructing readers/students to train or deploy a Huggingface model, hyperparameter tuning is often left out, and users end up with manual tracking on different training runs.
Figure 4 introduces a simple sample code to capitalize on the Automatic Model Tuning feature in Sagemaker and release users from performing “graduate student descent " (pinto2011). Alternatively, we encourage readers to explore the Syne Tune feature to perform more complex hyperparameter tuning tasks.
import sagemaker
from sagemaker.huggingface import HuggingFace
# gets role for executing training job
role = sagemaker.get_execution_role()
hyperparameters = {
’model_name_or_path’:’Helsinki-NLP/opus-mt-de-en’,
’output_dir’:’/opt/ml/model’,
’dataset_name’: ’wmt14’,
’source_lang’: ’de’,
’target_lang’: ’en’,
’dataset_config_name’: ’de-en’
# git configuration to download our fine-tuning script
git_config = {
’repo’: ’https://github.com/huggingface/transformers.git’,
’branch’: ’v4.36.0’}
# creates Hugging Face estimator
huggingface_estimator = HuggingFace(
entry_point=’run_translation.py’,
source_dir=’./examples/pytorch/translation’,
instance_type=’ml.g4dn.xlarge’,
instance_count=1,
role=role,
git_config=git_config,
transformers_version=’4.36.0’,
pytorch_version=’2.1.0’,
py_version=’py310’,
hyperparameters = hyperparameters,
# starting the train job
huggingface_estimator.fit()
Figure 4: Working Sample Code to Fine-tune a Translation model with Huggingface Transformers
import sagemaker
from sagemaker.huggingface import HuggingFace
from sagemaker.tuner import IntegerParameter
from sagemaker.tuner import HyperparameterTuner
# gets role for executing training job
role = sagemaker.get_execution_role()
hyperparameters = {
’model_name_or_path’:’Helsinki-NLP/opus-mt-de-en’,
’output_dir’:’/opt/ml/model’,
’dataset_name’: ’wmt14’,
’source_lang’: ’de’,
’target_lang’: ’en’,
’dataset_config_name’: ’de-en’,
metric_definitions=[
{"Name": "train_loss", "Regex": "’train_loss’: ([0-9]+(.|e\-)[0-9]+),?"}
# creates Hugging Face estimator
huggingface_estimator = HuggingFace(
entry_point=’my_run_translations.py’,
source_dir=’./scripts’,
instance_type=’ml.g4dn.xlarge’,
instance_count=1,
role=role,
transformers_version=’4.36.0’,
pytorch_version=’2.1.0’,
py_version=’py310’,
metric_definitions=metric_definitions,
hyperparameters = hyperparameters,
hyperparameter_ranges = {
"max_source_length": IntegerParameter(80,200),
"max_target_length": IntegerParameter(80,200),
objective_metric_name = "max_source_length"
objective_type = "Maximize"
hpo_definitons = [
{"Name": "max_source_length",
[…]
tuner = HyperparameterTuner(
huggingface_estimator,
objective_metric_name,
hyperparameter_ranges,
hpo_definitons,
max_jobs=10,
max_parallel_jobs=1,
objective_type=objective_type,
tuner.fit()
Figure 5: Working Sample Code to do Hyperparameter Tuning for a Translation model with Huggingface Transformers